OpenStudy (anonymous):

There is a clown's face on the top of a spinner. The tip of his hat rotates to (−2, 5) during one spin. What is the cosine value of this function?

OpenStudy (imstuck):



OpenStudy (imstuck):



OpenStudy (imstuck):

In order to find the cos of that angle, you need to find the value of x, since cos of an angle is the side opposite over the hypotenuse. Can you find x using Pythagorean's Theorem?

OpenStudy (anonymous):

I think, is it 5/-2?

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@Purplerainbowcherry

OpenStudy (anonymous):

I'll give a medal to whoever can help me solve this problem!!

OpenStudy (anonymous):

you need to use the pythagoras theorem to find x

OpenStudy (anonymous):

x^2 = 5^2 + (-2)^2

OpenStudy (anonymous):

x^2 = 25 + 4 x^2 = 29 x = 5.385

OpenStudy (anonymous):

ya exactly, now you can find the cos value

OpenStudy (anonymous):

opp/hyp?

OpenStudy (anonymous):

cos= adj/hyp which would be -2/5.383

OpenStudy (anonymous):

-0.3714

OpenStudy (anonymous):

i believe so

OpenStudy (anonymous):

these are my choices: -2 5 \[(5\sqrt{29})/29\] -2 sqrt 29 / 29

OpenStudy (anonymous):

the last one, as when you find it out you get the answer

OpenStudy (anonymous):

its d thanks somuch 1

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

although most of the credit goes to @IMStuck